OpenStudy (anonymous):

Differential Equation 4xy dx + y dx + 2ydy - 2x^2

OpenStudy (anonymous):

hi tyeyn haha

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

\[4xydx+ydx+2ydy-2x ^{2}dy=0 y(4x+1)dx+(xy-2x^{2})dy=0\]

OpenStudy (anonymous):

\[dx/dy-(2x^{2}/4x+1)(1/y)=-2/(4x+1)\]

OpenStudy (anonymous):

in original equation is it: ...-2x^2 or ...-2x^2 dy?

OpenStudy (anonymous):

if it's simple... - 2x^2 I think you should : 1) find solution of homogeneous equation (without 2x^2) 2) find particular solution (form: y=Ax^2 + Bx+C) 3) general solution: y=y(homogeneous)+y(particular) let me know if there is a problem

OpenStudy (anonymous):

inik, it's -2x^2 dy

OpenStudy (anonymous):

good! did you get the answer? i'm back on line... let me know if still a problem